From 50 life samples of mini drill bits made of low-carbon alloy steel, (expressed in the number of holes that can be drilled before it breaks), an average of 12.68 can be calculated, with a standard deviation of 6.83. Calculate a 95% confidence interval for the average mini drill in this condition.
1
Expert's answer
2020-05-04T19:47:21-0400

The Central Limit Theorem

Let "X_1,X_2,...,X_n" be a random sample from a distribution with mean "\mu" and variance "\sigma^2." Then if "n"

is sufficiently large, "\bar{X}" has approximately a normal distribution with "\mu_{\bar{X}}=\mu" and "\sigma_{\bar{X}}^2=\sigma^2/n."

If "n>30," the Central Limit Theorem can be used.

We need to construct the 95% confidence interval for the population mean "\mu."

The following information is provided: "\bar{X}=12.68, \sigma=6.83, n=50,"

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96." The corresponding confidence interval is computed as shown below:


"CI=\big(\bar{X}-z_c{\sigma\over\sqrt{n}},\bar{X}+z_c{\sigma\over\sqrt{n}}\big)="

"=\big(12.68-1.96\times{6.83\over\sqrt{50}},12.68+1.96\times{6.83\over\sqrt{50}}\big)\approx"

"\approx(10.787, 14.573)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "10.787<\mu<14.573," which indicates that we are 95% confident that the true population mean "\mu"

is contained by the interval "(10.787, 14.573)."